OpenStudy (anonymous):

please help me! stressd out! Write a system of two equations using x and y in each for the following. The sum of two numbers is 170 and their difference is 26. To skate at Roller Heaven, each person must pay a membership fee (m) and a fee for each session (f). Art attended five sessions and paid $25 in all. Later, when he had attended 10 sessions, he paid $40 in all. How much is the membership fee and how much are the sessions. Write a system of two equations with the variables m and f in each to represent this problem. To skate at Roller Heaven, each person must pay a membership fee (m)

OpenStudy (anonymous):

Simplify this expression: x2*x6What did you learn from that example? When the same variable is multiplied, any exponents are added together. Adding to the exponent is the same as multiplying more times: Rule 1: (xa)*(xb) = xa+b There are other rules with exponents as well. If multiply two variables adds their exponents, then division must subtract exponents! Check out this example: Well, remember that this is just a quick way of writing this: Hopefully you remember enough basic algebra that you know to cancel a factor that's in the numerator and the denominator. In fact, we can scratch off three of the x's, leaving just the numerator: So that gives us another rule: Rule 2: Let's introduce a few more rules of exponents quickly: Rule 3: x1 = x That rules makes sense, because having just one x can't equal anything else but x, right? The next one might make a little less sense, but here it is: Rule 4: x0 = 1. (ANYTHING to the 0 power = 1). Here's why this is the case. With exponents, we are always dealing with multiplication. The identity in multiplication is 1. Imagine taking rule 1, and adding in a couple 0's: (xa)*(xb) = xa+b+0+0 Logically, we shouldn't have changed anything by adding those zeros. Since you can rewrite it as: xa*xb*x0*x0 In order for that to be the same as xa*xb, the x0 factors must equal 1. Since x2 really just means x*x, and x6 just means 6 more x's, we end up with 8 x's multiplied together, right? Well what is an exponent -- the number of times we multiply something! Therefore, x2 times x6 equals x8.

OpenStudy (anonymous):

so here u go!!!!

OpenStudy (anonymous):

this is a EX: for u to still help u !!!! :)